Please help me with this problem? A virus takes up a volume of approximately 8.5x10-14 cubic centimeters. Calculate the estimated number of viruses in 28 cubic centimeters.
I tried to solve by o.ooooooooooooo85x28...and was not successful. Thank you

Answer by Fombitz(32388)

One virus takes up 8.5E-14 cc (cubic centimeters). N viruses take up 28 cc. Assuming that the viruses stack together and fill all of the space, set up a ratio.
Volume of one virus / One virus = Volume of N viruses / N viruses
Or using the information given,
8.5E-14 cc/1 virus = 28 cc/N (viruses)
Rearranging, by multiplying both sides by N and dividing by 8.5E-14, you get,
N = 28/8.5E-14 = 3.3E14 viruses.